OpenStudy (anonymous):

David can row 5 mph in still water. She takes as long to row 8 miles upstream as 23 downstream. How fast is the current?

ganeshie8 (ganeshie8):

\[\frac{8}{(5-x)} = \frac{23}{(5+x)} \] can you solve for x now ?

OpenStudy (anonymous):

I think so. Thank you!

ganeshie8 (ganeshie8):

great !! so u understand how we arrived at above equation ?